OpenStudy (anonymous):

Solve the system of equations. Y=2x^2-3 Y=3x-1 A. No solution B. (-1/2, 5), (2, -5/2) C. (-1/2, -5/2), (2, 5) D, (1/2, 5/2) , (2, 5)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=2x^2-3}\) \(\large\color{black}{ \displaystyle y=3x-1}\) in the first equation the 'y' is equivalent to '2x^2-3' and in the second equation (this same y' is equivalent to) '3x-1' so wouldn't that mean that 3x-1=2x^2-3 ?

OpenStudy (solomonzelman):

basically, given \(\large\color{black}{ \displaystyle a=b}\) and \(\large\color{black}{ \displaystyle a=c}\) therefore \(\large\color{blue}{ \displaystyle b=c}\).

OpenStudy (solomonzelman):

here, we go by the same exact principle, but the terms are not a,b,c.

OpenStudy (solomonzelman):

So, based on what I said, what do you think you should do?

OpenStudy (anonymous):

I'm so confused.. :(

OpenStudy (anonymous):

lol he basically blew up your brain! :'D I believe it's A no solution because n a linear equation and quadratic equation cannot come together :)

OpenStudy (anonymous):

That's what I was thinking but I just got really confused. Thank you.

OpenStudy (anonymous):

your welcome :)

OpenStudy (solomonzelman):

that is not true at all !!

OpenStudy (solomonzelman):

I bet 99.999999 % there is a solution at least 1.

OpenStudy (solomonzelman):

Anyway, do you understand the a=b a=c and therefore b=c principle that I posted ?

OpenStudy (anonymous):

Yeah. I read it over a couple of times.

OpenStudy (solomonzelman):

ok, so what you should then say is: \(\large\color{black}{ \displaystyle 2x^2-3=3x-1 }\)

OpenStudy (solomonzelman):

Then, solve this quadratic